# When Machine Unlearning Meets Retrieval-Augmented Generation (RAG): Keep Secret or Forget Knowledge?

### v1

@@ IV-A The Overview of RAG-based Unlearning
We propose a lightweight RAG-based unlearning framework for LLMs. This scheme relies on RAG technology, serving as a machine unlearning plugin for both closed-source and open-source LLMs. As shown in Figure 2 (b), the model owner only manipulates the RAG part. For the target data, he/she constructs its unlearned knowledge containing a confidentiality requirement. Given a prompt about the target data, the retriever can return the carefully designed knowledge. Then, a prompt template integrates the target data with the knowledge. In this case, the LLM will take into account the confidentiality requirement described in the knowledge, thus effectively achieving sample and concept unlearning. It is worth noting that the manual creation of the template is sufficient. In this work, the framework is detailed in Figure 4. It consists of three parts: instruction description, input prompt and retrieved knowledge. When an input prompt hits the knowledge base, the unlearned LLM must respond to the input based on the retrieved content; otherwise, must respond directly to the input.
@@ IV-B Identify Key Optimization
[…]
@@ IV-C Design of Unlearned Knowledge
The first component $Q$ ensures that unlearned knowledge can be retrieved. Otherwise, the LLM will respond directly to the original prompts, thus failing to maintain confidentiality. The second component $P$ is designed to restrict the LLM’s behavior. For a target $c$, our goal is to craft its unlearned knowledge $k$ that incorporates $P$ and $Q$, the process is described in Algorithm 1function Unlearned Knowledge Generation. When $k$ is used as the context of $c$, the LLM will provide a confidentiality response, such as refusing to answer. Obviously, optimizing the two components is challenging. We adopt a simple approach that concatenates the two non-overlapping components. Formally, $k=P+Q$, where ‘$+$’ denotes the textual concatenation. First, we craft $Q$ to induce the LLM to provide a refusal response, such as ‘Sorry, I cannot generate the related content due to copyright issues.’. Next, given $Q$, we further build $P$ to satisfy the retrieval condition. The unlearned knowledge encompasses both significant components, as illustrated in Figure 3.
@@ IV-C1 Crafting Constraint Component $Q$
When component $Q$ is used as context, the LLM will refuse to respond to the prompt $G(c)$ containing $c$. The key is to ensure that the LLM understands the confidentiality requirement embedded in $Q$, regardless of the prompt $G(c)$. A straightforward approach resembles prompt engineering, where $Q$ is constructed through optimization. However, there are three challenges with this approach. First, model owners may not have access to the parameters in LLM, especially in the case of closed-source models. Second, even in a white-box setting, where LLM access is possible, the computational resources are prohibitively high. Third, it remains uncertain whether an optimized $Q$ would be effective across different LLMs. To address these challenges, we employ a heuristic approach that does not rely on optimization. The detail is displayed in Algorithm 1function Craft Constraint.
@@
Here, $V$ specifies the length of $Q$ to conserve computational resources. In the function, using crafted $Q$ as context, $LLM_{un}$ is prompted to generate a response $R$ for a prompt related to $c$. If $R$ is still related to $c$, $Q$ is regenerated until $LLM_{un}$ refuses to respond or the maximum attempt (i.e., $T$ [^1] Note that if the maximum attempt is reached, $Q$ is the final crafted content.). In general, a single query is sufficient to craft a usable $Q$. Taking ‘Harry Potter’ as an example, the constraint component is as follows:
@@ IV-C2 Crafting Retrieval Component $P$
Target Sample. In this case, model owners have access to the trained samples $\mathbf{E}_{sample}=\{x_{i}|i=1,2,...,N\}$ that need to be forgotten. For each $x_{i}$, we assume $x_{i}$ has the highest similarity to itself. Therefore, we directly define $P=x_{i}$, resulting in $k=x_{i}+Q$. Although this design is straightforward, our experimental results demonstrate its high effectiveness in Section V-B.
Target Concept. In contrast to sample unlearning, model owners only have access to the concepts $\mathbf{E}_{concept}=\{C_{i}|i=1,2,...,N\}$ that need to be forgotten. For queries related to $C_{i}$, $P$ must comprehensively cover the knowledge related to $C_{i}$, thereby making $k$ easier to retrieve. In Equation 1, the retrieval function $\Gamma(\cdot)$ uses both semantic and keyword matching, similar to Blended RAG [36]. If model owners can access the retriever and input prompts, they will use optimized methods to generate $P$. However, in real-world scenarios, input prompts are often unavailable. Therefore, we use the unlearned model itself $LLM_{un}$ to produce thorough descriptions of $C_{i}$, improving the ranking of $P+Q$ during retrieval. For example, we employ the following prompt:
@@ IV-D Extend to MLLMs and LLM-based Agents
Closed Source GPT-4o N/A forgotten concept set
@@
llama-2-7b-chat-hf Tiny-nq forgotten concept set
@@ V-A Experimental Settings
Open-source LLMs. As shown in Table III, we used the Llama-2-7b-chat-hf model provided by Meta AI. Because the model is open-source, the trained data is accessible to model providers. In this scenario, our unlearning scheme can work for both sample and concept unlearning.
For sample unlearning, we fine-tuned a Llama-2-7b-chat-hf model with specific samples. Accordingly, we created a confidentiality requirement for each trained sample and modified the unlearned knowledge base. Unlike crafting questions for 100 topics, we applied a data extraction attack proposed by Carlini et al. [39] to construct a forgotten set. Specifically, we regarded the prefixes of trained samples as malicious prompts, inducing the LLM to regurgitate the suffixes of these samples. In the experiment, we selected a subset of the NQ dataset[18] that contains 2,000 natural question-and-answer pairs, known as Tiny-nq. Each item consists of a query, a long answer (i.e., a paragraph from a Wikipedia page), and a short answer (i.e., a specific phrase or entity). An example of the tiny-nq set is as follows:
@@
For concept unlearning, we randomly selected $100$ topics from Wikipedia and generated five related questions for each topic, thereby crafting a forgotten concept set. It is worth noting that the Llama-2-7b-chat-hf model should have learned these selected topics, i.e., generating related content. In this scenario, fine-tuning is unnecessary. We then created confidentiality requirements for these topics and updated the unlearned knowledge base. For both sample and concept unlearning, we used the forgotten set to query the llama-2-7b-chat-hf model, collected its responses to measure each unlearning scheme.
@@
Evaluation Measures. Different from Section II-B3, we evaluated the effectiveness of unlearning schemes using the data extraction attack. For each data point $x_{i}\in$ the forgotten set $\mathbf{E}$, we took the output of the original LLM (without unlearning) as its correct answer. During verification, we compared the output results before and after unlearning using the following two metrics. ROUGE measures the degree to which the generated output from the unlearned LLM matches the correct answer. We used the ROUGE-L recall score [40]. A low ROUGE score indicates that an unlearning scheme is effective. The unlearning success rate (USR) represents the probability of successful forgetting attempts. For each attempt, we applied an additional LLM (i.e., GPT-4o) to measure whether it achieved the unlearning goals. Specifically, we took the original input, forgotten target, and responses before and after unlearning as a new prompt and queried the GPT-4o model. In this case, a high USR demonstrates the effectiveness of an unlearning scheme. The template is as follows:
@@
To comprehensively evaluate these unlearning schemes, we also selected out-of-forget-scope sets containing normal prompts. This set evaluates the harmlessness of unlearning schemes, and we detail the process and measurement metrics in Section V-D. A well unlearning scheme should exhibit comparable performance to its original counterpart.
@@
ROUGE USR ROUGE USR ROUGE USR ROUGE USR
Concept GPT-4o N/A N/A 52.8 5.4% N/A N/A 0.0 100%
Gemini N/A N/A 71.9 3.0% N/A N/A 0.0 100%
llama-2-7b-chat-hf 61.3 32.4% 75.1 9.6% 82.4 39.0% 0.0 100%
Sample llama-2-7b-chat-hf 30.9 74.6% 73.2 21.7% 37.8 64.1% 0.0 100%
@@ V-B Effectiveness
In Section III-D, effectiveness is a key metric for evaluating unlearning schemes. Following the aforementioned effectiveness measurements, for each scheme, we reported its ROUGE and USR in open-source and closed-source LLMs, respectively.
@@
Closed-source LLMs. As shown in Figure I, gradient ascent and $\mu$-unlearning cannot be applied to closed-source models, and the related values are marked as N/A. Though in-context unlearning works well in polarity analysis tasks, it completely fails in the open Q&A task. We reported the effectiveness of two schemes in two closed-source models, as shown in Table IV.
@@
RAG-based unlearning relies on the retrieval accuracy of GPT-4o. For each prompt $x_{i}$ in the forgotten set, we recorded the hit success rate. Assume the related concept of $x_{i}$ is $C_{i}$, and the unlearned knowledge is $k_{i}$. Only if the retrieval model returns $k_{i}$ containing a confidentiality requirement from its external knowledge base, it is considered a successful hit. Notably, the advanced model can achieve a $100\%$ hit success rate due to its strong understanding ability, which is the intuition behind our scheme. For example, the hit knowledge for ‘Harry Potter’ is shown in Figure 3 (b), consisting of the correct information and confidentiality requirement. Based on the advanced retrieval model, RAG-based unlearning can fully ‘forget’ the target concepts on ChatGPT and Gemini platforms, achieving $100\%$ USR scores. Additionally, the ROUGE values indicate that the responses generated by our scheme deviate significantly from correct answers, which further demonstrates the effectiveness of our scheme. Clearly, the two closed-source models understand the confidentiality requirement in the post-crafted prompt, resulting in stronger forgetting effects.
Open-source LLMs. Using the aforementioned settings, we evaluated both sample and concept unlearning in the Llama-2-7b-chat-hf model. The experimental results are shown in Table IV. For sample unlearning, Figure 3 (a) shows the hit knowledge of a trained sample, consisting of the sample itself and a confidentiality requirement. In this case, RAG-based unlearning achieves a $100\%$ USR. Although the other three unlearning schemes can also make the LLMs forget trained samples, their USR scores are below $75\%$. In addition, these schemes require significant computational resources or cause unpredictable side effects. For concept unlearning, given a prompt from the forgotten set, the retrieval model also achieves highly accurate retrieval results. Similar to the closed-source models, it successfully achieves a $100\%$ USR on the forgotten set. In contrast, the three schemes can only reach USR scores below $40\%$.
@@
RAG-based Unlearning Against Harmful Output. Concept unlearning can be applied to a scenario involving the removal of harmful output. Due to the variation in regulations across different regions, deploying the same LLM across regions requires different safety guardrails. Our scheme enables the customization of LLMs to meet regional regulations and impose restrictions on specific harmful output. In this case, we conducted a concept unlearning experiment on some harmful topics. Specifically, we randomly selected 25 malicious topics, such as murder, robbery, and cybercrime. For each of these topics, we used an additional LLM (i.e., GPT-4o) to create five related questions, constructing a forgotten set. It is worth noting that closed-source LLMs have established safety guardrails against these harmful topics. Therefore, we generated ten question-and-answer pairs [^3] For each topic, the question-and-answer pairs do not overlap with the forgotten set. for each harmful topic to craft a training set. The Llama-2-7b-chat-hf model was fine-tuned on the training set and exhibited a $91.2\%$ probability of generating harmful output on the forgotten set. We then evaluated four unlearning schemes on the fine-tuned model. For the other three schemes, they treat the training samples as forgotten targets, thereby implementing concept unlearning. In addition to the USR metric, the harmful output probability (HOP) is used in place of the ROUGE metric, as tracking the reduction in harmful generation provides a more accurate measure in this scenario.
@@
Gradient Ascent 43.2% 30.4%
In-context Unlearning 76.2% 8.4%
$\mu$-Unlearning 41.6% 36.8%
RAG-based Unlearning 0.8% 99.2%
@@
In summary, the existing schemes are not robust against strict concept unlearning and can only forget the collected samples. However, in theory, any advanced retrieval model can perform the precise retrieval process, as discussed in Section VII-A, thereby fully ‘forgetting’ the target samples or concepts.
@@
-chat-hf
Gradient Ascent w/o Rephrase N/A N/A 32.4%
w/ Rephrase N/A N/A 26.6%
In-context Unlearning w/o Rephrase 5.4% 3.0% 9.6%
w/ Rephrase 5.6% 2.6% 5.8%
$\mu$-Unlearning w/o Rephrase N/A N/A 39.0%
w/ Rephrase N/A N/A 17.8%
RAG-based Unlearning w/o Rephrase 100% 100% 100%
w/ Rephrase 100% 100% 97.4%

### v2

@@ IV-A The Overview of RAG-based Unlearning
We propose a lightweight RAG-based unlearning framework for LLMs. This scheme relies on RAG technology, serving as a behavioral unlearning plugin for both closed-source and open-source LLMs. As shown in Figure 2 (b), the model provider only manipulates the RAG part. For the target data, she constructs its unlearned knowledge containing a confidentiality requirement. Given a prompt about the target data, the retriever can return the carefully designed knowledge. Then, a prompt template integrates the target data with the knowledge. In this case, the LLM will take into account the confidentiality requirement described in the knowledge, thus effectively achieving sample and concept unlearning. It is worth noting that the manual creation of the template is sufficient. In this work, the framework is detailed in Figure 4. It consists of three parts: instruction description, input prompt and retrieved knowledge. When an input prompt hits the knowledge base, the unlearned LLM must respond to the input based on the retrieved content; otherwise, must respond directly to the input.
In practical scenarios, we consider a more challenging threat model in which an adversarial user knows the RAG-based unlearning framework. Leveraging this knowledge, the adversary crafts adaptive prompts to bypass the system’s confidentiality guarantees. We detailed this threat model in Section VII-A and provided a thorough analysis of its potential impact on the effectiveness of our scheme.
@@ IV-B Identify Key Optimization
where $Verify(\cdot)$ is a discriminator that can determine whether $A(;\theta)$ has forgotten $C_{i}$, $G(\cdot)$ is a generator that can create related questions for each concept, and $\Gamma(G(C_{i});\textbf{UK}\cup\textbf{BK})$ is a retriever that can extract related knowledge $K_{i}$ of $G(C_{i})$ from $\textbf{UK}\cup\textbf{BK}$. To forget Econcept, the optimized function focuses on the unlearned knowledge set UK, achieving a large value in Equation 1.
@@ IV-C Design of Unlearned Knowledge
The first component $Q$ ensures that unlearned knowledge can be retrieved. Otherwise, the LLM will respond directly to the original prompts, thus failing to maintain confidentiality. The second component $P$ is designed to restrict the LLM’s behavior. For a target $c$, our goal is to craft its unlearned knowledge $k$ that incorporates $P$ and $Q$, the process is described in Algorithm 1function Unlearned Knowledge Generation. When $k$ is used as the context of $c$, the LLM will provide a confidentiality response, such as refusing to answer. Obviously, optimizing the two components is challenging. We adopt a simple approach that concatenates the two non-overlapping components. Formally, $k=P+Q$, where ‘++’ denotes the textual concatenation. First, we craft $Q$ to induce the LLM to provide a refusal response, such as ‘Sorry, I cannot generate the related content due to copyright issues.’. Next, given $Q$, we further build $P$ to satisfy the retrieval condition. The unlearned knowledge encompasses both significant components, as illustrated in Figure 3.
@@ IV-C1 Crafting Constraint Component $Q$
When component $Q$ is used as context, the LLM will refuse to respond to the prompt $G(c)$ containing $c$. The key is to ensure that the LLM understands the confidentiality requirement embedded in $Q$, regardless of the prompt $G(c)$. A straightforward approach resembles prompt engineering, where $Q$ is constructed through optimization. However, there are three challenges with this approach. First, model providers may not have access to the parameters in LLM, especially in the case of closed-source models. Second, even in a white-box setting, where LLM access is possible, the computational resources are prohibitively high. Third, it remains uncertain whether an optimized $Q$ would be effective across different LLMs. To address these challenges, we employ a heuristic approach that does not rely on optimization. The detail is displayed in Algorithm 1function Craft Constraint.
@@
Here, $V$ specifies the length of $Q$ to conserve computational resources. In the function, using crafted $Q$ as context, $LLM_{un}$ is prompted to generate a response $R$ for a prompt related to $c$. If $R$ is still related to $c$, $Q$ is regenerated until $LLM_{un}$ refuses to respond or the maximum attempt (i.e., TT [^1] Note that if the maximum attempt is reached, $Q$ is the final crafted content.). In general, a single query is sufficient to craft a usable $Q$. Taking ‘Harry Potter’ as an example, the constraint component is as follows:
@@ IV-C2 Crafting Retrieval Component $P$
Target Sample. In this case, model providers have access to the trained samples $\mathbf{E}_{sample}=\{x_{i}|i=1,2,...,N\}$ that need to be forgotten. For each $x_{i}$, we assume $x_{i}$ has the highest similarity to itself. Therefore, we directly define $P=x_{i}$, resulting in $k=x_{i}+Q$. Although this design is straightforward, our experimental results demonstrate its high effectiveness in Section V-B.
Target Concept. In contrast to sample unlearning, model providers only have access to the concepts $\mathbf{E}_{concept}=\{C_{i}|i=1,2,...,N\}$ that need to be forgotten. For queries related to $C_{i}$, $P$ must comprehensively cover the knowledge related to $C_{i}$, thereby making $k$ easier to retrieve. In Equation 1, the retrieval function $\Gamma(\cdot)$ uses both semantic and keyword matching, similar to Blended RAG [42]. If model providers can access the retriever and input prompts, they will use optimized methods to generate $P$. However, in real-world scenarios, input prompts are often unavailable. Therefore, we use the unlearned model itself $LLM_{un}$ to produce thorough descriptions of $C_{i}$, improving the ranking of $P+Q$ during retrieval. For example, we employ the following prompt:
@@ IV-D Extend to MLLMs and LLM-based Agents
Closed
Source
GPT-4o N/A forgotten concept set
@@
llama-2-7b-chat Tiny-nq forgotten concept set
@@ V-A Experimental Settings
Open-source LLMs. As shown in Table III, we used the Llama-2-7b-chat model provided by Meta AI. Because the model is open-source, the trained data is accessible to model providers. In this scenario, our unlearning scheme can work for both sample and concept unlearning.
For sample unlearning, we fine-tuned a Llama-2-7b-chat model with specific samples. Accordingly, we created a confidentiality requirement for each trained sample and modified the unlearned knowledge base. Unlike crafting questions for 100 topics, we applied a data extraction attack proposed by Carlini et al. [34] to construct a forgotten set. Specifically, we regarded the prefixes of trained samples as malicious prompts, inducing the LLM to regurgitate the suffixes of these samples. In the experiment, we selected a subset of the NQ dataset[21] that contains 2,000 natural question-and-answer pairs, known as Tiny-nq. Each item consists of a query, a long answer (i.e., a paragraph from a Wikipedia page), and a short answer (i.e., a specific phrase or entity). An example of the tiny-nq set is as follows:
@@
For concept unlearning, we randomly selected $100$ topics from Wikipedia and generated five related questions for each topic, thereby crafting a forgotten concept set. It is worth noting that the Llama-2-7b-chat model should have learned these selected topics, i.e., generating related content. In this scenario, fine-tuning is unnecessary. We then created confidentiality requirements for these topics and updated the unlearned knowledge base. For both sample and concept unlearning, we used the forgotten set to query the llama-2-7b-chat model, collected its responses to measure each unlearning scheme.
@@
Evaluation Metrics. We evaluated the effectiveness of unlearning schemes using the data extraction attack. For each data point $x_{i}\in$ the forgotten set $\mathbf{E}$, we took the output of the original LLM (without unlearning) as its correct answer. During verification, we compared the output results before and after unlearning using the following two metrics.
  • ROUGE. It measures the degree to which the generated output from the unlearned LLM matches the correct answer. We used the ROUGE-L recall score [45], which computes the length of the longest common subsequence between the generated text and the reference answer. A lower ROUGE-L score reflects a larger divergence from the original reference, indicating that the unlearning scheme is more effective.
  • Unlearning Success Rate (USR). It represents the proportion of successful forgetting attempts out of all attempts. For each attempt, we used an external LLM (i.e., GPT-4o) to assess whether the unlearning goals are achieved. In this setting, a more effective unlearning scheme produces more successful attempts, resulting in a higher USR score.
Specifically, we constructed a prompt using the original input, the forgotten target, and the model’s responses before and after unlearning, and queried GPT-4o for evaluation. The used template is as follows:
@@
Following prior studies [11, 35], we employed the MIA to quantify residual memorization of target samples or concepts after unlearning. We benchmarked several representative MIAs (LOSS [33], ZLib [34], GradNorm [35] and Min-K [36]) on LLMs, and selected Min-K, the best-performing one, for the final evaluation. Specifically, given a text, Min-K computes the log-likelihood of each token under the target LLM, sorts tokens by this value, retains the lowest $k\%$ (the hardest tokens), and uses their mean log-likelihood as the test statistic. Prior empirical analyses suggest $K=20$ works best, and we adopt this setting in our experiments. For each evaluation, the forgetting set served as the member set, and news articles from 2025 as the non-member set. The metric is as follows:
  • TPR@1%FPR. It is the true positive rate at a fixed 1% false positive rate, used as the metric of residual memorization. On the original LLM (without unlearning), Min-K achieved 4.1%. If the LLM has completely forgotten the target data, its performance should approach that of random guessing, i.e., 1%. Thus, a TPR@1%FPR closer to 1% indicates less residual memorization and, consequently, a more effective unlearning scheme.
To ensure the results’ credibility, each experiment was conducted five times and reported as mean $\pm$ standard deviation.
@@
ROUGE USR TPR@1%FPR ROUGE USR TPR@1%FPR ROUGE USR TPR@1%FPR ROUGE USR TPR@1%FPR
[…]
@@ V-B Effectiveness
For each scheme, we reported its ROUGE, USR and TPR@1%FPR in open-source and closed-source LLMs.
@@
Closed-source LLMs. As shown in Figure I, gradient ascent and $\mu$-unlearning cannot be applied to closed-source models, and the related values are marked as N/A. Though in-context unlearning works well in polarity analysis tasks, it completely fails in the open Q&A task. In Table IV, we reported the effectiveness of two schemes in two closed-source models.
@@
RAG-based unlearning relies on the retrieval accuracy of GPT-4o. For each prompt $x_{i}$ in the forgotten set, we recorded the hit success rate. Assume the related concept of $x_{i}$ is $C_{i}$, and the unlearned knowledge is $k_{i}$. Only if the retrieval model returns $k_{i}$ containing a confidentiality requirement from its external knowledge base, it is considered a successful hit. Notably, the advanced model can achieve a $100\%$ hit success rate due to its strong understanding ability, which is the intuition behind our scheme. For example, the hit knowledge for ‘Harry Potter’ is shown in Figure 3 (b), consisting of the correct information and confidentiality requirement. Based on the advanced retrieval model, RAG-based unlearning can fully ‘forget’ the target concepts on ChatGPT and Gemini platforms, achieving an average USR of $99.3\%$. The low average ROUGE score of 0.04 reflects a substantial deviation from correct answers, and the TPR@1%FPR of 1.4% is close to the ideal 1%, further confirming the scheme’s effectiveness. Clearly, the two closed-source models understand the confidentiality requirement in the post-crafted prompt, resulting in stronger forgetting effects. In comparison, the baseline produces a USR of 5.2%, a ROUGE of 61.7, and a TPR@1%FPR of 3.2%. Relative to these values, RAG-based unlearning under closed-source LLMs achieves significantly higher USR, and lower ROUGE and TPR@1%FPR, highlighting its superior unlearning performance.
Open-source LLMs. Using the aforementioned settings, we evaluated both sample and concept unlearning in the Llama-2-7b-chat model. The experimental results are shown in Table IV. For sample unlearning, Figure 3 (a) shows the hit knowledge of a trained sample, consisting of the sample itself and a confidentiality requirement. All three baselines enable LLMs to forget target samples to some extent, with gradient ascent performing best, with a USR of 75.8%, a ROUGE score of 32.7 and a TPR@1%FPR of 2%. In contrast, our RAG-based unlearning achieves perfect forgetting (100% USR), maximum deviation from the correct answers (0.0 ROUGE), and a low TPR@1%FPR of 1.2%, indicating better unlearning performance. For concept unlearning, given prompts from the forgotten set, the retrieval model still achieves highly accurate retrievals. Consistent with results on closed-source LLMs, RAG-based unlearning attains a 99.8% USR on the forgotten set, whereas all three baselines remain below 40%. The low TPR@1%FPR of 1.3% further confirms that our scheme effectively suppresses residual memorization.
@@
RAG-based Unlearning Against Harmful Output. Concept unlearning can be applied to a scenario involving the removal of harmful output. Due to the variation in regulations across different regions, deploying the same LLM across regions requires different safety guardrails. Our scheme enables the customization of LLMs to meet regional regulations and impose restrictions on specific harmful output. In this case, we conducted a concept unlearning experiment on some harmful topics. Specifically, we randomly selected 25 malicious topics, such as murder, robbery, and cybercrime. For each of these topics, we used an additional LLM (i.e., GPT-4o) to create five related questions, constructing a forgotten set. It is worth noting that closed-source LLMs have established safety guardrails against these harmful topics. Therefore, we generated ten question-and-answer pairs [^3] For each topic, the question-and-answer pairs do not overlap with the forgotten set. for each harmful topic to craft a training set. The Llama-2-7b-chat model was fine-tuned on the training set and exhibited a $91.2\%$ probability of generating harmful output on the forgotten set. We then evaluated four unlearning schemes on the fine-tuned model. For the other three schemes, they treat the training samples as forgotten targets, thereby implementing concept unlearning. In addition to the USR metric, the harmful output probability (HOP) is used in place of the ROUGE metric, as tracking the reduction in harmful generation provides a more accurate measure in this scenario.
@@
[…]
@@
In summary, the existing schemes are not robust against strict concept unlearning and can only forget the collected samples. However, in theory, any advanced retrieval model can perform the precise retrieval process, as discussed in Section VII-C, thereby fully ‘forgetting’ the target samples or concepts.
@@
-chat
Gradient
Ascent
[…]
In-context
Unlearning
[…]
RAG-based
Unlearning
[…]